tantheta/1-cos^2theta
1
Expert's answer
2013-03-01T05:48:51-0500
tantheta/(1-cos^2theta)=(sin(theta)/cos(theta))/sin^2(theta)=1/( cos( theta ) * sin( theta ) )=2/( 2sin(theta)*cos(theta)
)=2/sin(2*theta)=2csc(2theta)